PROBING THE STRUCTURE OF IONIC LIQUID SURFACES BY ROTATIONALLY AND ELECTRONI-CALLY INELASTIC SCATTERING OF NO

\underline{M. P. ZIEMKIEWICZ}, A. ZUTZ, and D. J. NESBITT , JILA, University of Colorado and National Institute of Standards and Technology, Boulder, Colorado, USA.

Room temperature ionic liquids (RTILs) are a highly diverse class of materials with many potential technological applications. They are candidates for use in advanced electrolytes, green solvents, and supported liquid membranes for CO_{2} sequestration. We present studies where inelastic scattering of high or low velocity nitric oxide provides insight into the microscopic structure of these complex surfaces. As an open shell diatomic, jet-cooled NO [^{2}Π_{1/2}(J = 0.5)] features both molecular and electronic collision dynamics as seen by probing scattered rotational and spin-orbit distributions respectively. These studies show substantial variation in degree of rotational and electronic excitation as ionic liquid identity is varied. Also, surface heating is found to have a strong effect on scattered spin-orbit branching, possibly due to the dependence of surface structure on temperature. This is discussed in terms of a picture where the electronic degree of freedom may serve as a sensitive measure of the cationic versus anionic nature of the top few layers of this material.